2 of 6 people are exposed to a safety hazard in a minivan. All of them are from different states. Texas, Alabama, Mississippi, Louisiana, Iowa, Illinois. What is the probability that either Texan or Iowan are exposed to the hazard?
Answer #1

The probability that a person randomly picked from this 6 is exposed to a safety hazard is p=\frac{2}{6}=\frac{1}{3}

That means the probability that Texan is exposed is 1/3. The probability that Iowan is exposed is 1/3

P(\text{Texan})= P(\text{Iowan})=\frac{1}{3}

the probability that either Texan or Iowan are exposed to the hazard is

P(Texan U lowan) = P(Texan) + P(Iowan)-P(Texan n Iowan) but the event of Texan being exposed is independent of Iowan being ex

ans: the probability that either Texan or Iowan are exposed to the hazard is 0.5556